OpenStudy (anonymous):

In quadrilateral PQRS, the coordinates are P(0, 0), Q(a + c, 0), R(2a + c, b), and S(a, b). How can you use coordinate geometry to show that the diagonals are perpendicular? A. Apply the Distance Formula to show that opposite sides line PQ and line RS are congruent. B. Find the slopes of line PQ and line RS , and show that their product is 1. C. Apply the Distance Formula to show that opposite sides line PS and and line QR are congruent. D. Find the slopes of line PR and line QS , and show that their product is -1.

hartnn (hartnn):

so, do you know distance formula and slope formula ?

hartnn (hartnn):

Distance between points (x1,y1) and (x2,y2) is \(\huge d=\sqrt{(x_1-x_2)^2+(y_1-y_2)^2}\) The slope of the line through points (x1,y1) and (x2,y2) is given by : \(\huge m=\frac{y_1-y_2}{x_1-x_2}\)

OpenStudy (anonymous):

I dont understand this ?

hartnn (hartnn):

"Apply the Distance Formula to show that opposite sides line PQ and line RS are congruent." so we need to find PQ and RS and you know co-ordinates of all 4. so try to apply the distance formula and find PQ first ?

hartnn (hartnn):

the same way you did for your previous prolem

OpenStudy (anonymous):

PQ = sqrt (0-a+c)^2+(0-0)^2 ?? like this ?

hartnn (hartnn):

[0 - (a+c) ]^2 be careful with brackets

hartnn (hartnn):

\(PQ = \sqrt{[0-(a+c)]^2+0} = a+c\)

OpenStudy (anonymous):

ok but what's the answer am gonna get for this equation ?

hartnn (hartnn):

so thats PQ try to find RS now ?

OpenStudy (anonymous):

ok

hartnn (hartnn):

then in same way find PS and QR too for part C

OpenStudy (anonymous):

RS = sqrt [ (2a+c) - a ] ^2 + [ b-b ] ^2

hartnn (hartnn):

good, simplify that...

OpenStudy (anonymous):

and how do I do that ?

hartnn (hartnn):

(2a+c) - a = 2a-a +c = ... ?

OpenStudy (anonymous):

i got 0 ? they're kinda the same

hartnn (hartnn):

didn't you get a+c for both PQ and RS ? which is what we wanted to prove , that PQ= RS

hartnn (hartnn):

now find PS And QR using same distance formula, and check whether they come out to be same

OpenStudy (anonymous):

yea sorry am not good at this thing so I don't know how to the things you're telling me

hartnn (hartnn):

actually, its the same way you found PQ you did it on your own. whats the problem in finding PS and QR in same way ?

OpenStudy (anonymous):

is it C ?

OpenStudy (anonymous):

or A

hartnn (hartnn):

we're done with A, now we're trying to solve part C

OpenStudy (anonymous):

so A is not the answer so we're gonna repeat everything we did before but with PS and QR

hartnn (hartnn):

ok, lets use slope formula. find slope between PQ and RS first.

hartnn (hartnn):

P(0, 0), Q(a + c, 0) so, \( \large m= \dfrac{0-0}{a+c-0} =... ? \)

OpenStudy (anonymous):

\[\frac{ 0 }{ a+c }\]

hartnn (hartnn):

which is 0 so the product won't be -1! right ?? for lines to be perpendicular, the product of slopes should be -1 and since its not option B, which option remains ?

OpenStudy (anonymous):

if the slope should be -1 then its D

hartnn (hartnn):

the 'product' of slopes should be -1 and yes, its D :)

OpenStudy (anonymous):

ok thanks a lot

hartnn (hartnn):

welcome :)

OpenStudy (anonymous):

can you help me with this http://openstudy.com/study#/updates/533bb2fee4b06e93c844f393